Bernie Corporation has 200,000 shares of $10 par common stock outstanding and 100,000 shares of $100 par, 6% cumulative, nonparticipating preferred stock outstanding. Dividends have not been paid for the past two years. This year, a $3,000,000 dividend will be paid. What are the dividends per share payable to preferred and common, respectively?

a. $6; $12. b. $18; $6.

c. $6; $6. d. None of these answer choices is correct
Answer #1

Calculate dividend per share :

Annual dividend preferred stock = 100000*100*6% = 600000

Preferred dividend = 600000*3 = 1800000/100000 = 18 per share

Common stock dividend = (3000000-1800000)=1200000/200000 = 6 per share

so answer is b) 18;6